# Ting En Lam1,2, Yuhan Chen1,3, Elston Tan1,4, Eric Peh1, Ruirui Chen1, Paritosh Parmar1, Basura Fernando1

### v1

@@ 3.2 Unique Characteristics of our dataset
    Multi-level answers. VideoQA datasets generally provide single-level answers. Single-level answers might contain less information. Multi-level answers are valuable because they offer a richer and more nuanced perspective, which leads to deeper insights—considering not only the immediate circumstances but also underlying psychological, social, and personal factors that contributed to characters’ decisions. By employing multi-level answers alongside thoughtful questions, we can cultivate deeper reasoning and analysis of characters’ actions, enabling the training and evaluation of models on complex question-answer pairs that demand a comprehensive understanding of longer causal chains. Further elaborated discussion on the importance of multi-level answers to ‘Why’-questions included in Supplementary material.
@@
    Grounded in diverse Visual & Motion information. The Tom and Jerry cartoon series is visually rich and dynamic with intricate scenes, actions, and character interactions varying across all episodes. We asked annotators to craft questions where the answer is grounded in the video. By grounding questions(-answers) in the visual input, the model is required to analyze the video to gain access to a broader range of details and cues/clues from the videos to provide meaningful answers. For example, the same question—Why was Jerry laughing?—appears in two different seasons (episodes S03E17, S05E13). However, the Explanations are different: 1) for S03E17, the Explanation is “Jerry found it funny that the seal stole the fish from Tom."; 2) while for S05E13, the Explanation is “Jerry was laughing at Tom who fell down the stairs and into the fountain." This concretely shows that VideoQA models need to understand visual information to correctly answer the questions. Grounding QAE set in vision is underscored by the presence of unusual situations in cartoons—typically not found in real-world cases. For example, Jerry removing Tom’s nose and fitting strawberry in place of Tom’s nose or Lion hushing Jerry. Due to this, models cannot answer just based on the correlations/biases in its training data, rather models ought to process and understand the contents thoroughly to answer correctly. This is crucial even in with large language models (LLMs)/embeddings, which are also susceptible to such frequency-based biases/hallucinations [20, 24, 60]. For example, Lion is closely related to killing in embeddings from LLMs; in a qualitative analysis, we found that models may tend to choose an answer containing ‘killing’ Jerry as the motive for Lion-character’s action, while in actuality the Lion-character was trying to quiet Jerry. Revealing such biases has been considered in other computer vision problems [4, 12], but is yet to be introduced in VideoQA. Our dataset offers a valuable opportunity to address this challenge. Additionally, understanding Tom & Jerry cartoons requires modeling of details and actions at varying levels of granularity—from sweeping movements to subtle emotional cues via facial expressions.
@@
    Focus on causal reasoning in visually dynamic scenes. Our dataset focuses on causal reasoning in Tom and Jerry cartoons, which have an abundance of dynamic and rapidly changing scenes, making it challenging for traditional video understanding models to track and comprehend the context effectively. To effectively model cause-and-effect relationships and answer causal questions, video understanding models must comprehend the temporal and contextual flow of events, linking them together amid a multitude of rapid scene changes and dynamic interactions. Furthermore, characters and objects may appear, disappear, and reappear, potentially limiting the model’s access to only partial observations. We term this as Dynamic Scene Linking (DSL). DSL is different from temporal modeling in traditional sense, which is typically concerned with modeling temporal evolution of events within a scene (more details in Supplementary material). Related challenges have been recognized in other computer vision problems [34, 45], but not yet in VideoQA. We hypothesize that humans effectively understand these cartoons by forming a mental ‘world model’ of the scene. This mental model aids in bridging gaps between discontinuous scenes and provides a coherent representation of the overall context. Incorporating such capabilities into VideoQA has the potential to enhance model performance. Notably, current VideoQA datasets often lack coverage of such complex scenarios, unlike our dataset.
@@
    More challenging and requiring cognitive effort. Understanding causal chains can help explain why things happen and how different factors influence each other. Identifying and solving longer causal chains in the process of answering causal (why) questions involve linking together multiple cues/clues, which requires significant cognitive efforts on the part of VideoQA models, especially, when the clues/cues are embedded in different scenes (described above) like in our dataset. Annotators were explicitly instructed to craft QAE pairs that require linking together multiple cues under multiple scene changes. It is difficult to quantify the complexity of causal questions. To the best of our knowledge, there are no existing approaches to accomplish this. To this end, we propose using Causal Chain Length as a way to measure the complexity of causal questions. Computing lengths of Causal Chains. We leverage GPT-4 [32] to compute the lengths of causal chains of 100 randomly chosen QAE pairs from our dataset and existing causal(-Why) VideoQA datasets [47, 17, 19]. Results are presented & compared in Figure 2. We observed that the QAE pairs in our dataset have significantly longer causal chains than existing QA datasets. This finding sheds light on the unique challenges posed by our dataset and underscores the importance of developing robust VideoQA models capable of handling complex causal reasoning tasks. Examples of causal chains and additional details on computing causal chain lengths provided in Supplementary material.
@@
[…]
    [loop,autoplay,poster=1,width=0.95]3Figs/tor_frs/fr_0015
    Figure 3: Examples of various types of reasoning required by our dataset. Please zoom in and view in AdobeReader to watch the embedded videos for better explanation.
@@
    Wide spectrum of types of reasoning required. Our dataset demands various types of reasoning like 1) Deductive; 2) Inductive; 3) Spatial; 4) Causal; 5) Critical thinking; 6) Emotion; 7) Abductive; and 8) Temporal. We have visualized examples demonstrating these reasoning types in Figure 3. It is also possible that a question may involve more than one type of reasoning. We conducted human studies to determine the types of reasoning required by our dataset and NextQA and observed that our dataset demands a wider range of reasoning than NextQA (results in 2(a)). More information and definitions of types of reasoning and human studies are provided in Supplementary material.
  7. 7.
    Character names used instead of general nouns. In our dataset, we employ character names rather than generic nouns, enhancing contextual understanding and knowledge representation. The utilization of character names necessitates character/actor resolution, encouraging Video QA models to better comprehend question context and character dynamics. This contextual understanding results in more precise and contextually fitting answers. This approach becomes particularly crucial in scenes with multiple instances of the same type of animals, such as several cats or mice. Relying on character names ensures that models are required to distinguish between individual characters and accurately identify who is interacting with whom. Failure to discern these interactions may lead to inaccurate responses.
@@ 3.3 Benchmarks
Our dataset comprises of following types of questions.
@@
    Multiple Choice Question Answering. VideoQA models are provided with the question, the associated video footage, the correct answer, and four incorrect answers; the task for the models is to pick the correct answer.
@@
    Open-ended Answer Generation. The VideoQA models are provided only with the question and the video footage, and the models have to generate the answers in natural language.
Incorrect answer/option mining.
In designing MCQA, the answer options should be sufficiently distinct from both the correct answer & each other. However, they should also possess enough semantic similarities to challenge reasoning capabilities beyond common sense. This way models cannot answer correctly by taking shortcuts like simple object/action matching in vision-language spaces. Naive Negative mining or using randomly chosen answers from other questions as incorrect options is unproductive; it renders the correct answer too obvious, so we avoid this strategy. Rather, we introduce the following strategies.
@@
    Vanilla Hard Negative mining. A pre-trained Sentence-Bert was adopted in the measure of semantic similarity. For each question of interest, we first sample top-10 other questions based on the cosine-similarity measure of the Sentence-Bert embedding. From these 10 question candidates, the associated answers are then sampled, and the cosine-similarity measure is once again applied between the correct answer and the answer candidate. Using a similarity measure in the equation: $\alpha cossim(Q_{i},Q_{n})\cdot\beta cossim(A_{i},A_{n})$, where $\alpha$ and $\beta$ can be varied to adjust the weights coming from the questions or answers, the top 4 answers are included as the option candidates. To prevent candidates from being too similar or rephrasing the form of the answer, the measure heavily penalizes answers with a high similarity $cossim(A_{i},A_{n})$ by setting $\beta$ = 0.001 to reduce the overall score of the candidate.
@@
    Causal-Confusion Negative mining. We leverage the power of large language models to generate confusing incorrect options which have the opposite effect (e.g. Tom was hitting Jerry $\rightarrow$ Tom was not hitting Jerry) or the causal agents are swapped (e.g. Tom was hitting Jerry $\rightarrow$ Jerry was hitting Tom). These causally-opposite incorrect options play an important role in causal action question-answering datasets like ours. We hypothesize that the models might not be doing sufficient joint modeling of language and vision. As such, Video QA models might be able to answer based on matching actors and objects in questions and answers. However, when tested on Causally-Opposite incorrect options, simple question-answer matching might not be enough, and consequently, Video QA models may fail to answer correctly. We hypothesize that simple objects and actors might be modeled by VideoQA models, but relationships might not be as they require more dynamic features and as such are harder to model. We have provided examples from Causal-Confusion vs. Vanilla Hard Negative in the Supplementary material.
Design of train-val-test splits.
In general, we divide the dataset into a train-val-test split ratio of 70:15:15. We also use the following train/test strategies.
  1. 1.
    Seen-episode testing.
    1. (a)
      Uniformly Distributed Seen-episode testing (UD). In this case, dataset samples are randomly divided into train-val-test splits. No constraints such as the train samples should be drawn only from certain episodes or only from certain parts of certain episodes are imposed. Due to this, it is likely that uniformly distributed chunks/portions of all episodes from all seasons are seen during training. This way the VideoQA models can potentially have sparse storylines with them.
    2. (b)
      Consecutive Partially Seen-episode testing (PS). In this case, train set is constrained to be drawn from starting 70% video of each episode; val set is constrained to be drawn from the middle 70-85% video of each episode; while the test set is constrained to be drawn from the remaining last 15% of the video of each episode. This creates a scenario where the model has partially “seen" episodes during training stage and then does “educated guessing" when answering during testing stage.
  2. 2.
@@ 4 Experiments
We benchmark the performance of state-of-the-art VideoQA methods on our newly introduced causal action QA dataset on two tasks: 1) MCQA; 2) OEAG.
@@
BlindQA [2] 29.38 13.07 26.51 11.54 25.13 11.02
EVQA [2] 29.38 13.48 31.32 13.32 27.82 14.65
CoMem [8] 32.08 13.88 26.10 09.89 23.12 09.27
HME [6] 32.35 14.02 29.53 12.36 25.13 10.22
HCRN [16] 32.48 16.98 32.01 14.97 25.67 12.23
HGA [15] 31.40 15.36 29.40 13.19 28.23 13.84
MIST [7] 60.51 43.67 59.89 38.05 53.36 36.29
@@
BLIP-2 [18] 43.67 23.32 45.88 24.18 46.64 26.48
Video-LLaMA [57] 35.00 11.73 35.16 9.34 29.62 9.68
VideoChat2 [21] 38.14 15.36 38.91 15.93 40.83 15.99
Table 1: MCQA Results on our dataset. A: answer; E: explanation.
4.1 Benchmarking on MCQA
Performance metric.
Accuracy is used as the performance metric.
@@ 4.1 Benchmarking on MCQA
Following prior work [47], we benchmark the performance of the following models on our causal action QA dataset: BlindQA [2]; EVQA [2]; CoMem [8]; HME [6]; HCRN [16]; HGA [15]. Additionally, we also evaluate recent vision-language models like MIST [7], BLIP-2 [18] and multimodal instruction tuning models Video-LLaMA [57], VideoChat2 [21], which obtain solid results on vision-language tasks including VideoQA. We have provided further details on each of these models in Supplementary material.
Evaluation Protocols.
Models are evaluated under the following protocols.
  • Protocol 1. Models’ ability to select the correct Answer is only evaluated. Their ability to select the correct explanation is not taken into consideration. The chance accuracy in this protocol is $1/5$. Results from this protocol are presented under columns marked as A in Table 1.
  • Protocol 2. Models’ ability to select the correct Answer as well as the correct Explanation is evaluated. If the model can select the correct answer, but not the correct explanation, then the model is considered to have failed. In this protocol, the chance accuracy becomes $1/5\times 1/5=1/25$. Results from this protocol are presented under columns marked as A+E in Table 1.
The performances of the baseline models for both protocols are presented in Table 1. In general, we observe that the performances of most of the baseline approaches are low. Most of the models gain only around 11-12% improvement over the chance accuracy for both protocols. MIST seems to be doing exceptionally well compared to other models. We hypothesize that this could be because MIST being geared for long-form video understanding, is better able to handle long video contexts, which are frequent in our dataset. Nonetheless, overall, we observed the following failure modes: 1) Limited Evidence Consideration: They often focus on a small subset of evidence, neglecting other relevant cues distributed throughout the sequence, which is problematic for datasets with long causal chains like ours. 2) Models failed because they tried to exploit shortcuts like object-noun or action-verb matching in video and language spaces instead of focusing on causal relationship modeling—such shortcuts can lead to wrong predictions on our dataset due to our hard negative mining, correct and incorrect answers would likely contain these nouns/verbs. The most effective way to discriminate on our dataset is by inferring the causal relationship. We have discussed further shortcomings and provided qualitative results in Supplementary material.
Comparing the two protocols, we find that Protocol 2, where the model has to select both the correct answer and the correct explanation is significantly more difficult than selecting just the correct answer for all the models. Comparing splits, we observe that, in line with our intuition, UD seems to be the easiest split across all the models. UD is followed by PS; with UN being the most difficult split. This order of difficulty is intuitive. It can also indicate that having at least some understanding of past events or storylines may aid in reasoning about current events or actions.
Negative type Accuracy (%)
Vanilla Hard 60.51
Causal-Confusion 40.57 ($\downarrow 20\%$)
Table 2: Impact of causally confusing options.
We further conducted an experiment where we tested the best performing model, MIST, on our Causal-Confusion set, where incorrect answer choices have the same objects and actors as the correct answer option, but cause and effect relationships are reversed or altered. We observed a significant drop in performance as shown in Table 2. This could potentially be due to causal relationships being not modeled adequately even by such advanced VideoQA models. Interestingly, we also found the S-BERT similarity score to be above 90% for an action (e.g., Tom is running after Jerry) and its Causal-Confusion version (e.g., Tom is not running after Jerry); while these sentences would be opposite/different in terms of human perception. We believe that Causality might be more overlooked than we think in various fields, not just in computer vision.
Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
A E A E A E A E A E A E A E A E
BlindQA [2] 0.2412 0.1912 0.1039 0.0618 0.0415 0.0221 0.1243 0.0891 0.2732 0.1885 0.2956 0.0703 0.1034 0.0403 0.4987 0.4195 2.7646
UATT [51] 0.2661 0.2203 0.1257 0.0805 0.0575 0.0330 0.1409 0.1046 0.3171 0.2293 0.3547 0.2775 0.1197 0.0397 0.5495 0.4767 3.3928
HME [6] 0.2650 0.2072 0.1301 0.0713 0.0554 0.0219 0.1504 0.0914 0.3125 0.2273 0.3184 0.0826 0.0936 0.0639 0.5328 0.4737 3.0975
HGA [15] 0.2891 0.2263 0.1588 0.0897 0.0789 0.0323 0.1856 0.1130 0.3388 0.2399 0.3980 0.2426 0.2782 0.0573 0.6026 0.4561 3.7872
BlindGPT-2 0.3770 0.3165 0.2530 0.1749 0.1632 0.1091 0.2452 0.1761 0.3858 0.2966 0.4305 0.3529 1.2482 0.7690 0.6755 0.6271 6.6006
VisionGPT-2 0.3878 0.3095 0.2605 0.1727 0.1738 0.1091 0.2560 0.1756 0.3934 0.2941 0.4498 0.3385 1.3725 0.7539 0.6760 0.6350 6.7582
BLIP-2 [18] 0.1381 0.0815 0.0451 0.0256 0.0167 0.0059 0.0664 0.0480 0.1618 0.1312 0.0530 0.0422 0.2279 0.1046 0.3837 0.3614 1.8931
Video-LLaMA [57] 0.1241 0.1181 0.0419 0.0344 0.0115 0.0098 0.1477 0.1163 0.1719 0.1435 0.2055 0.1383 0.0836 0.0430 0.5734 0.4834 2.4464
VideoChat2 [21] 0.2353 0.2116 0.0823 0.0776 0.0250 0.0253 0.1769 0.1295 0.2667 0.2168 0.3264 0.2547 0.3980 0.2910 0.6445 0.5908 3.9524
(a) (a)
Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
A E A E A E A E A E A E A E A E
BlindQA [2] 0.2193 0.1795 0.0772 0.0468 0.0326 0.0149 0.1122 0.0813 0.2501 0.1767 0.2960 0.1147 0.1099 0.0434 0.5129 0.4260 2.6935
UATT [51] 0.2693 0.1947 0.1257 0.0581 0.0440 0.0146 0.1466 0.0847 0.3226 0.2107 0.3255 0.2012 0.1444 0.0646 0.5364 0.4829 3.2260
HME [6] 0.2475 0.1830 0.1031 0.0549 0.0363 0.0154 0.1417 0.0845 0.2820 0.2173 0.2933 0.2549 0.0831 0.0655 0.5165 0.4929 3.0719
HGA [15] 0.2586 0.1842 0.1085 0.0517 0.0365 0.0148 0.1433 0.0932 0.2909 0.1806 0.2908 0.2128 0.0877 0.0266 0.5231 0.4078 2.9111
(b) (b)
Model BLEU-1 BLEU-2 BLEU-3 METEOR ROUGE SPICE CIDEr S-BERT CapsMIX
A E A E A E A E A E A E A E A E
BlindQA [2] 0.2312 0.1847 0.0869 0.0550 0.0319 0.0198 0.1263 0.0930 0.2519 0.1806 0.2758 0.2447 0.0667 0.0265 0.4844 0.4153 2.7747
UATT [51] 0.2659 0.1936 0.1180 0.0645 0.0548 0.0223 0.1423 0.0868 0.3025 0.2200 0.3449 0.2595 0.0940 0.0713 0.5447 0.4750 3.2601
HME [6] 0.2640 0.1930 0.1238 0.0644 0.0556 0.0239 0.1457 0.0868 0.2998 0.2155 0.3115 0.1623 0.0944 0.0698 0.5341 0.4761 3.1207
HGA [15] 0.2674 0.2293 0.1337 0.0823 0.0620 0.0308 0.1703 0.1123 0.3186 0.2363 0.3734 0.2558 0.1972 0.0477 0.5765 0.4692 3.5628
(c) (c)
Table 3: OEAG Results on our dataset. (a) UD split; (b) PS split; (c) UN split.
4.2 Benchmarking on OEAG
CapsMIX Performance metric.
We measure the performance of open-ended generation in terms of BLEU-1,2,3 [33], METEOR [3], ROUGE [23], SPICE [1], CIDEr [44] and Sentence-BERT [36] scores by comparing with the ground truth answers as done by captioning and QA literature. However, we note that with such a wide range of metrics, it is difficult to get a comprehensive insight into models’ performances and compare them. To address that, we introduce a comprehensive metric, termed Caps-MIX (Captioning Metrics Integration eXpert), which integrates all the previously mentioned scores after normalizing them to their theoretical best values. This not only makes it more accessible to compare models based on just a single number but also combines the characteristics of individual metrics, each of which measures performance from a unique perspective. We note that the WUPS [27] score was originally developed for measuring the performance of single-word answers, & may not be an ideal metric for our dataset, which contains longer, detailed answers. Thus, we refrain from using it.
Baselines.
@@ 4.2 Benchmarking on OEAG
Following prior work [47], we benchmarked the performance of EVQA [2]; UATT [51]; HME [6]; HGA [15] on open-ended answer generation task. Additionally, we also report the zero-shot performance of recent multimodal video understanding models BLIP-2 [18], Video-LLaMA [57], VideoChat2 [21].
We evaluated the baselines for generating answers and explanations. The performances of various models are presented in Table 3. Comparing the BLEU scores, we can see that the performances degrade as we move from BLEU-1 to BLEU-3, indicating that the models are struggling with generating longer answers. We confirmed this qualitatively. Comparing split-type wise, we found that as in the case of MCQA, UD was the easiest split. Overall, we conclude that models struggle quite a bit with open-ended generation.
@@
Inspired by the success of large language models (LLMs), we experimented with leveraging GPT-2 [35], a publicly available LLM, as our natural language answer generator. We found that a pre-trained version out of the box did not work well—it generated mostly random, unrelated words. However, upon simple training on the train set, it performed significantly better than the baselines we considered as shown in Table 3(a) (BlindGPT-2). However, it is unclear if these large language models have “seen" Tom and Jerry scripts during their pretraining stage. If so, then pretraining on relevant scripts followed by finetuning on our dataset could potentially be a reason for GPT-2’s good performance. Although this is less likely. What is more likely to be the reason behind this good performance is the language modeling/generating capability of GPT-2.

### v2

@@ 3.2 Unique Characteristics of our CausalChaos! Dataset
    Multi-level answers offer a richer and more nuanced perspective, which leads to deeper insights—considering not only the immediate circumstances but also underlying psychological, social, and personal factors that contributed to characters’ decisions. By employing multi-level answers alongside thought-provoking questions, we can cultivate deeper reasoning and analysis of characters’ actions, enabling the training and evaluation of models on complex question-answer pairs that demand a comprehensive understanding of longer causal chains. Further elaborated discussion on the importance of multi-level answers to ‘Why’-questions included in Appendix.
@@
    Grounded in diverse Visual & Motion information. The Tom & Jerry cartoon series is visually rich and dynamic, with intricate scenes, actions, and character interactions varying across episodes. We had annotators craft questions where the answer is grounded in the video. This grounding compels the model to analyze the video for a broader range of details and clues to provide meaningful answers. For example, consider Figure 4(a), same question appears in different episodes, but their answers/context are completely different. This demonstrates that it is crucial for VideoQA models to understand visual information to answer correctly on our dataset. Multimodal nature of VideoQA task is enhanced by the presence of unusual situations in cartoons—typically not found in real-world datasets as shown in Figure 4(b,c). Since the answers are embedded in videos, models cannot rely on correlations/biases in their training data; they must thoroughly process and understand the content to answer correctly. This is crucial even with large language models (LLMs) and embeddings, which are also susceptible to frequency-based biases and hallucinations [20, 24, 60]. For example, Lion is often associated with killing in LLM embeddings. In a qualitative analysis, we found that models might choose an answer involving ’killing’ Jerry as the motive for the Lion-character’s action, while the Lion-character was actually trying to quiet Jerry Figure 4(b-left). Revealing such biases has been considered in other computer vision problems [4, 12], but is yet to be introduced in VideoQA. Our dataset offers a valuable opportunity to address this challenge. Additionally, understanding Tom & Jerry cartoons requires modeling of details and actions at varying levels of granularity—from sweeping movements to subtle emotional cues via facial expressions.
@@
    Focus on causal reasoning in visually dynamic scenes. Our dataset focuses on causal reasoning in Tom and Jerry cartoons, characterized by dynamic and rapidly changing scenes. Characters and objects may appear, disappear, and reappear, limiting the model’s access to partial observations. Video understanding models struggle to track and understand context. To answer causal questions, these models must link events to form a causal chain despite rapid scene changes and dynamic interactions. We term this Dynamic Scene Linking (DSL), distinct from traditional temporal modeling, which typically focuses on the gradual transitions within a scene as shown in Figure 5 (more details in Appendix). Related challenges are noted in other computer vision problems [34, 45] but not yet in VideoQA. We hypothesize that humans understand cartoons by forming a mental “world model" of the scene, which helps bridge gaps between discontinuous scenes and provides a coherent overall context. Current VideoQA datasets often lack coverage of such complex scenarios, unlike our dataset.
@@
    More challenging and requiring cognitive effort. We used GPT-4o [31] to compute the lengths of causal chains of 100 randomly chosen QAE pairs from our dataset and existing causal(-Why) VideoQA datasets [47, 17, 19]. Results are presented in Figure 3. We observed that the QAE pairs in our dataset have longer causal chains than existing QA datasets, potentially suggesting our QAs are more complex. Identifying and solving longer causal chains in the process of answering causal (why) questions involve linking together multiple cues/clues, which requires significant cognitive efforts on the part of VideoQA models, especially, when the clues/cues are embedded in different scenes (described above), as in our dataset.
@@
    Figure 6: Examples of various types of reasoning required by our dataset. Please zoom in & view in AdobeReader to play the embedded videos. Further types of reasoning visualized in Appendix.
@@
    Wide spectrum of types of reasoning required. Our dataset demands various types of reasoning like 1) Deductive; 2) Inductive; 3) Spatial; 4) Causal; 5) Critical thinking; 6) Emotion; 7) Abductive; and 8) Temporal. We have visualized examples demonstrating these reasoning types in Figure 6. It is also possible that a question may involve more than one type of reasoning. We conducted human studies to determine the types of reasoning required by our dataset and NextQA and observed that our dataset demands a wider range of reasoning than NextQA (results in 3(a)). More information and definitions of types of reasoning and human studies are provided in Appendix.
@@ 3.3 Benchmarks
Our dataset comprises two types of questions. The first is Multiple Choice Question Answering (MCQA). VideoQA models are provided with the question, the associated video footage, the correct answer, and four incorrect answers; the task for the models is to pick the correct answer. The second is Open-ended Answer Generation (OEAG). The models are provided only with the question and the video footage, and the models have to generate the answers in natural language.
Incorrect answer/option mining.
In designing MCQA, answer options should be distinct from the correct answer & each other but semantically similar enough to challenge reasoning beyond commonsense. This prevents models from taking shortcuts like simple object/action matching in vision-language spaces. Using randomly chosen answers as incorrect options is unproductive and makes the correct answer too obvious, so we avoid this strategy. Instead, we introduce these strategies:
@@
    Vanilla Hard Negative mining. A pre-trained Sentence-Bert measures semantic similarity. For each question, we sample the top-10 other questions based on cosine similarity of Sentence-Bert embeddings. From these, we sample associated answers and again apply cosine similarity between the correct and candidate answers. Using the equation $\alpha cossim(Q_{i},Q_{n})\cdot\beta cossim(A_{i},A_{n})$, with adjustable weights $\alpha$ and $\beta$, we select the top 4 answer candidates. To avoid overly similar answers, we penalize high similarity scores by setting $\beta$ to 0.001, reducing the candidate’s overall score.
@@
[…]
Design of Train-Val-Test splits.
We typically partition the dataset into 70% for training, and 15% each for validation and testing. Additionally, we employ specific strategies for training and testing.
@@
    Uniformly Distributed Seen-episode testing (UD). Dataset samples are randomly divided into train-val-test splits without constraints on which episodes or parts are included in each split. Consequently, uniformly distributed chunks of all episodes from all seasons are likely seen during training, potentially giving VideoQA models sparse storylines.
@@
    Consecutive Partially Seen-episode testing (PS). The train set is drawn from the first 70% of each episode, the val set from the middle 70-85%, & the test set from the last 15%. This setup allows the model to have partially "seen" episodes during training & make "educated guesses" during testing.
  3. 3.
@@ 4 Experiments
We benchmark SOTA VideoQA methods on our new causal action QA dataset for two tasks: 1) MCQA and 2) OEAG. Then, we explore if our dataset aids real-world QA.
4.1 Benchmarking on MCQA
Baselines.
Following prior work [47], we benchmark the performance of these models on our causal action QA dataset: BlindQA [2], EVQA [2], CoMem [8], HME [6], HCRN [16], and HGA [15]. We also evaluate recent vision-language models like MIST [7], BLIP-2 [18], and multimodal instruction tuning models Video-LLaMA [57] and VideoChat2 [21], GPT-4o [31] which excel on vision-language tasks. Further details on each model are provided in Appendix.
Evaluation Protocols.
Models are evaluated under these protocols, with Accuracy as the metric:
  • Protocol 1. Models’ ability to select the correct Answer (A) is only evaluated. Their ability to select the correct explanation is not taken into consideration. The chance accuracy in this protocol is $1/5$. Results from this protocol are presented under columns marked as A in Table 1.
  • Protocol 2. Models’ ability to select the correct Answer (A) as well as the correct Explanation (E) is evaluated. If the model can select the correct answer, but not the correct explanation, then the model is considered to have failed. In this protocol, the chance accuracy becomes $1/5\times 1/5=1/25$. Results from this protocol are presented under columns marked as A+E in Table 1.
@@ 4.1 Benchmarking on MCQA
BlindQA 29.38 13.07 26.51 11.54 25.13 11.02
EVQA 29.38 13.48 31.32 13.32 27.82 14.65
CoMem 32.08 13.88 26.10 09.89 23.12 09.27
HME 32.35 14.02 29.53 12.36 25.13 10.22
HCRN 32.48 16.98 32.01 14.97 25.67 12.23
HGA 31.40 15.36 29.40 13.19 28.23 13.84
MIST 62.22 44.88 62.22 42.86 55.91 37.90
MIST-CC 63.34 46.80 62.50 43.54 56.18 39.25
@@
BLIP-2 43.67 23.32 45.88 24.18 46.64 26.48
Video-LLaMA 35.00 11.73 35.16 9.34 29.62 9.68
VideoChat2 38.14 15.36 38.91 15.93 40.83 15.99
GPT-4o 63.64 48.17 63.84 49.79 64.82 52.23
Table 1: MCQA Results on our dataset.We design multitask version of MIST that learns to generate causal chains as an auxiliary task.
Quantitative results.
The performances of the baseline models for both protocols are presented in Table 1. In general, we observe that the performances of most of the baseline approaches are low. Most of the models gain only around 11-12% improvement over the chance accuracy for both protocols. MIST seems to be doing exceptionally well compared to other models. We hypothesize that this could be because MIST being geared for long-form video understanding, is better able to handle long video contexts, which are frequent in our dataset. Overall, we observed the following failure modes for models: 1) Limited Evidence Consideration: Models often focus on a small subset of evidence, neglecting other relevant cues distributed throughout the sequence, which is problematic for datasets with long causal chains like ours. 2) Models failed because they tried to exploit shortcuts like object-noun or action-verb matching in video-language spaces instead of focusing on causal relationship modeling—such shortcuts can lead to wrong predictions on our dataset due to our hard negative mining, correct & incorrect answers would likely contain these nouns/verbs. The most effective way to discriminate on our dataset is by inferring the causal relationship. We have discussed further shortcomings with qualitative results in Appendix. To mitigate these shortcomings, we design a MIST-based multitask model, MIST-CC, which, in addition to doing VideoQA, also learns to generate the causal chain from Video-Question features (groundtruth causal chains are generated by GPT-4 from QAE pairs). We found that this simple modification mildly boosts the performance. We have provided further details on MIST-CC in Appendix. Moving to VLM/MLLMs, we observe that they perform better than traditional models ([2, 15, 16]), except MIST. However, GPT-4o released only in late May 2024, outperformed all the models. Although, it is closed source model, we hypothesize its vision and language capabilities maybe significantly better than other VLMs. It was able to incorporate small details like facial expressions. To glean insights into it, we also asked it to give us its reasoning, and found that it analyzes each option individually and then selects the most likely answer. This is close to (or at least mimicking) how humans would approach this task. We believe this, at least on surface, seems to be going beyond just correlation-based answer picking as in non-VLM/MLLMs.
Comparing the two protocols, we find that Protocol 2, where the model has to select both the correct answer and the correct explanation is significantly more difficult than selecting just the correct answer for all the models. Comparing splits, We observe that UD is the easiest split across all models, followed by PS, with UN being the most difficult. This intuitive order suggests that understanding past events or storylines may aid in reasoning about current events or actions.
Negative type MIST MIST-CC GPT-4o
Vanilla Hard 62.22 63.34 63.64
Causal-Confusion 55.80 58.76 54.95
Table 2: Impact of causal-confusion.
We further conducted an experiment where we tested the best performing models, MIST, MIST-CC, & GPT-4o, on our Causal-Confusion set, where incorrect answer choices have the same objects and actors as the correct answer option, but cause-and-effect relationships are reversed or altered. We observed a significant drop in performance of all models as shown in Table 2. This could potentially be due to causal relationships being not modeled adequately even by such advanced VideoQA models. Performance of MIST-CC dropped relatively less, potentially because the auxiliary task of generating causal chains may have enhanced the understanding of causal relationships. Interestingly, we also found the S-BERT similarity score to be above 90% for an action (e.g., Tom is running after Jerry) and its Causal-Confusion version (e.g., Tom is not running after Jerry); while these sentences would be opposite/different in terms of human perception. We believe that Causality might be more overlooked than we think in various fields, not just in computer vision.
4.2 Benchmarking on Open-ended Answer Generation (OEAG)
CapsMIX Performance metric.
We measure the performance of OEAG in terms of BLEU-1,2,3 [33], METEOR [3], ROUGE [23], SPICE [1], CIDEr [44] & Sentence-BERT [36] scores by comparing with the ground truth answers as done by captioning & QA literature. The wide range of metrics complicates model comparison, so we introduce Caps-MIX (Captioning Metrics Integration eXpert), which normalizes and integrates all scores into a single metric, simplifying comparisons and combining the unique strengths of individual metrics.
Model CapsMIX
BlindQA 2.7646
UATT 3.3928
HME 3.0975
HGA 3.7872
BlindGPT-2 6.6006
VisionGPT-2 6.7582
BLIP-2 1.8931
Video-LLaMA 2.4464
VideoChat2 3.9524
GPT-4o 2.9851
Table 3: OEAG results on our dataset (UD split).
@@ 4.2 Benchmarking on Open-ended Answer Generation (OEAG)
Following prior work [47], we benchmarked the performance of EVQA [2]; UATT [51]; HME [6]; HGA [15] on OEAG task. We also report the zeroshot performance of recent multimodal video understanding models BLIP-2 [18], Video-LLaMA [57], VideoChat2 [21], GPT-4o [31].
We evaluated the baselines for generating answers and explanations. The performances of various models are summarized in Table 3. For full results, see Appendix. Overall, we observe that models struggle significantly with open-ended generation, including some recent VLMs and MLLMs. Despite GPT-4o’s strong performance on MCQA, it also falters on OEAG. These models likely perform better on MCQA by eliminating incorrect choices, but they fail to genuinely understand videos and perform causal reasoning for standalone answer generation. VideoChat2 seem to be doing exceptionally well, perhaps, because it was specifically trained on various video understanding tasks and datasets, including causal reasoning task. Comparing split-type wise, we found that as in the case of MCQA, UD was the easiest split.
@@
Inspired by the success of LLMs, we experimented with leveraging GPT-2 [35], a publicly available LLM, as our natural language answer generator. We found that a pre-trained version out of the box did not work well—it generated mostly random, unrelated words. However, upon simple training on the train set, it performed significantly better than the baselines we considered as shown in Table 3 (BlindGPT-2). However, it is unclear if these LLMs have “seen" Tom and Jerry scripts during their pretraining stage. If so, then pretraining on relevant scripts followed by finetuning on our dataset could potentially be a reason for GPT-2’s good performance. Although this is less likely. What is more likely to be the reason behind this good performance is the language modeling/generating capability of GPT-2.